(x^2D^2-xD+1)y=xlogx


1
Expert's answer
2021-03-12T06:54:04-0500

"x^2 y''-xy'+y=x\ln x"

This equation is linear, but non-homogeneous. So, at the first step we seek a particular solution of the homogeneous equation "x^2 y''-xy'+y=0".

We seek a solution of the form "y=x^{\alpha}" . Then

"0=x^2 y''-xy'+y=x^2 \alpha(\alpha-1)x^{\alpha-2} - x\alpha x^{\alpha-1} +x^{\alpha} =(\alpha-1)^2x^{\alpha}"

If "\alpha=1" then we have a solution "y=cx" , with "c" be an arbitrary constant.


The second step is variation of constants. We replace the constant "c" with a function "u(x)". Then

"y=xu"

"y'=xu'+u"

"y''=xu''+2u'"

and the origin equation transforms to the following one:

"x^2(xu''+2u')-x(xu'+u)+xu=x\ln x"

"x^3u''+x^2u'=x\ln x"

"xu''+u'=x^{-1}\ln x"

"(xu')'=x^{-1}\ln x"

"xu'=\int \ln x \frac{dx}{x}=\frac{1}{2}\ln^2x+c_1"

"u'=\frac{1}{2}x^{-1}\ln^2x+c_1x^{-1}"

"u=\int( \frac{1}{2}x^{-1}\ln^2x+c_1x^{-1})dx=\int(\frac{1}{2}\ln^2x+c_1)d\ln x=\frac{1}{6}\ln^3 x+c_1\ln x+c_2"

"y=xu=\frac{1}{6}x\ln^3 x+c_1x\ln x+c_2x"

This is a general solution of ODE.